find an equation of the line having the give slope and containing the give point m=-7, (3,4)
Answer by ankor@dixie-net.com(15747)

find an equation of the line having the give slope and containing the give point m=-7, (3,4)
:
Use point/slope formula: y - y1 = m(x - x1)
y - 4 = -7(x - 3)
y - 4 = -7x + 21
y = -7x + 21 + 4
y = -7x + 24